OpenStudy (anonymous):

Find a polynomial P(x) of degree 3 whose zeros are -3, 1, and 2.

OpenStudy (anonymous):

\[ P(x) = (x+3)(x-1)(x-2) = x^{3}-3x ^{2}-4x+6\]

OpenStudy (anonymous):

P(x)=(x+3)(x−1)(x−2)=x3−3x2−4x+6